If the test stat, t = 2.10 then we fail to reject the null that the population correlation coefficient is equal to zero (n = 6, alpha = 0.05).

True or false
Answer #1

Given, t=2.10 n-6, 1-6, L-0105 degrees of 0-2 = 6-2=4 freedom Pa value = 20 6 T4 <2.10) = 0.10 Since p-value 0.10 > 0.05, it